OpenStudy (anonymous):

Solve: x2 – x = 2 a) x = -1, x = -2 b) x = -2, x = 1 c) x = 1, x = 2 d) x = 2, x = -1

OpenStudy (anonymous):

d) x=2, x=-1

OpenStudy (anonymous):

put x= 2 u get 2^2 - 2 = 4-2 = 2 put x = -1, u get (-1)^2 - (-1) = 1 + 1 = 2 if happy with the answer, pls click GOOD ANSWER button :)